OpenStudy (anonymous):

13/18 divided by 5/6

myininaya (myininaya):

just flip second number and multiply

myininaya (myininaya):

\[\frac{13}{18} \cdot \frac{6}{5}\]

myininaya (myininaya):

6/18 can be reduced just so you know

myininaya (myininaya):

1/3=6/18 so we can write \[\frac{13}{18} \cdot \frac{6}{5}=\frac{13}{6} \cdot \frac{1}{5}\]

myininaya (myininaya):

now multiply straight across on top and straight across on bottom

OpenStudy (anonymous):

NO WAY!!! this is the answer 13/18 * 6/5= 13/3*1/5= 13/15(ANSWER)

myininaya (myininaya):

oops ues that 6 was a type-o suppose to be 3

myininaya (myininaya):

remember 6/18=1/3 not 1/6